You collected the following data in lab:

                          VHCl = 50.15 mL         CHCl = 0.960 M

                          VNaOH = 48.90 mL     CNaOH = 1.10 M

                          Tinitial ,HCl = Tinitial ,NaOH = Tinitial calorimeter = 25.7°C

                          Tf = 31.8°C

Assume that:

  • SNaOH = sHCl = swater = 4.18 J/g-°C
  • Ccal = 20.5 J/oC
  • Solutions can be treated as dilute (i.e., physical properties such as density are approximately that of water)
  1. Show one set of typed calculations for the following values: msoln, Tsoln, qsoln, qneut, mole of the limiting reactant, and molar Hneut. Express molar Hneut in KJ/mol.

Can you show your work with an explanation so ill understand?
Answer #1

सध्या का आतिरिक्त कुछ fore write on this margin n = VⓇC limiting reactant NOH = Shu Iwater = 4.18 J 7- VHe= 50-15 ml Cha=0.96

excess NaoH moles = 0.05379 -0.048 m = 5:079x103 moles. meras = 5.79 x 10 roles x 40 g 10.23169] since Speerific heat is same